OpenStudy (anonymous):

The product of a positive number and that number decreased by 9 is 52. Write and solve an equation that models the sentence above.

OpenStudy (whpalmer4):

a positive number: \(x\) that number decreased by 9: \(x-9\) product of a positive number and that number decreased by 9: \( x*(x-9)\) product " " " " is 52:\( x(x-9) = 52\) Solve \[x(x-9) = 52\]for \(x\). Make sure your solution is positive!

OpenStudy (anonymous):

any ideas

OpenStudy (anonymous):

ok so how would you solve that?

OpenStudy (whpalmer4):

You could factor it, you could complete the square, you could use the quadratic formula, you could graph it...

OpenStudy (whpalmer4):

Let's start by expanding and rearranging it a bit: What do you get if you use the distributive property to get rid of the ( ), and move all of the terms to one side of the = sign?

OpenStudy (anonymous):

you would get x^2-9=52

OpenStudy (anonymous):

if you further simplify it it would be x^2=61

OpenStudy (whpalmer4):

Uh, no, that would be wrong :-) x(x-9) =

OpenStudy (anonymous):

ohh ok then thanks

OpenStudy (whpalmer4):

\[x(x-9)=\] ???

OpenStudy (anonymous):

im confused :(

OpenStudy (whpalmer4):

\[x(x-9) = x*x -9*x =\]

OpenStudy (anonymous):

x^2-9x

OpenStudy (whpalmer4):

That's better. So after we move the 52 to the other side, we have \[x^2-9x-52 = 0\]Which of the methods that I listed do you know?